# Speculative Diffusion Decoding: Accelerating Language Generation Through Diffusion

### v3

@@ 5.2 Experimental Setup
Evaluation is conducted on three standard natural language processing tasks: (1) text summarization using the CNN/DM dataset [21], (2) text generation on the OpenWebText (OWT) dataset [22], and (3) text generation using MT Bench [43]. In each setting the model is queried for 1024 tokens using a greedy decoding scheme (temperature = 0). For the experiments, we evaluate the pretrained target models $M_{p}$ GPT-2 XL (1.5B), GPT-NEO (2.7B), and Vicuna (33B). We use Masked Diffusion Language Model (110M) as our drafter model $M_{q}$, which is a comparable size to the baseline drafter GPT-2 (86M) employed for our standard speculative decoding baseline [18].
@@
Our method is assessed empirically by walltime speedup and acceptance rate $\alpha$. The reported results are compared to recognized baselines of vanilla autoregressive decoding, standard speculative decoding (SpS) implementations as proposed by Leviathan et al., Chen et al., which our method most closely resembles. Additionally, we provide comprehensive comparison to these method, as well as the current fastest-to-date speculative decoding approaches, detailing the improvements SpecDiff provides in reduction of FLOPs and memory footprint in Table 1.
@@ 5.3 Results and Discussion
Table 3: Evaluation on MT Bench using SpecDiff (Ours), EAGLE, EAGLE-2, and SpS with Vicuna target models. FLOPs/step computes that floating point operations for the drafter to generate each sequence. ‘Overall Speedups’ for the baselines are measured on all Spec-Bench [36] tasks. denotes results collected using Spec-Bench.
@@
To provide comparison to other state-of-the-art speculative decoding methods, evaluation is conducted on generation prompts from MT Bench [43] using Vicuna 33B as the target model. Generation lengths are reduced to 512 tokens to better align with the task. For this setting, MDLM (141M) is pretrained with the Llama 1 tokenizer for compatibility with the target model used. As Vicuna is instruction tuned, we utilize a teacher-student instruction tuning approach resembling the knowledge distillation approach proposed by Zhou et al. to align the output distribution of MDLM with the target model. This approach is similar to the additional instruction tuning conducted to align Eagle and Eagle-2 on ShareGPT instruction tuning dataset [41, 42]. Because of this finetuning step, it is not necessary to use sequence initialization as with pretrained models. Notably, for task-specific generation on MT Bench, SpecDiff provides roughly $85\%$ of the speedup of Eagle-2, a state-of-the-art method. In terms of overall generation speed, it offers near-identical performance while just slightly edging out the top spot. Meanwhile, SpecDiff requires $\mathit{33\%}$ fewer floating point operations per draft generation. Thorough comparison to Eagle-2 and other speculative decoding strategies is highlighted in Table 1.
@@
While previous implementations of speculative decoding rely on a common architecture between the drafter and target models [12, 24], using smaller versions of the same architecture to generate draft sequences, these experiments demonstrate a robustness to using a completely different architecture for sequence drafting. This is particularly significant given the absence of finetuning in the reported results (Tables 2 and 4). Pretrained diffusion models can be directly purposed as draft models requiring no additional training.
@@ 6 Limitations and Future Work
We note that available discrete diffusion models are based on the GPT-2 tokenizer, which would restrict one immediate compatibility with target models using different tokenization schemes. For our experiments, we indeed trained a new discrete diffusion model from scratch with a different tokenizer, a process that demanded substantial computational resources and time. All our models will be released on HuggingFace and thus be directly used by the community.
@@
7 Conclusion
@@ Appendix A Drafter Convergence Analysis
In prior studies, theoretical results have focused on determining the optimal $\gamma$ to maximize the throughput of speculative decoding methods [12]. For SpecDiff, extending $\gamma$ introduces minimal overhead and becomes less important to consider. The number of sequential diffusion operations, $T$, instead impacts $\alpha$ as increasing this number improves the convergence of $M_{q}(x)$ to the learned distribution, aims to closely approximate the distribution of $M_{p}(x)$. Hence, an implicit dependency arises between $T$ and $\alpha$, which is reflected in Figure 2 (right).
@@
We are now ready to connect this to Equation 2:
Appendix B Accepted Draft Lengths
Figure 3: Accepted draft lengths for OpenWebText evaluation (left and middle) and MT Bench evaluation (right).
We notice a starch contrast in the lengths of accepted drafts between these two settings. While the distribution of accepted draft lengths in Figure 3 (right) is what would be anticipated given the lower values of $\alpha$, the longer generations on OpenWebText (left and middle of Figure 3) speak to the parallelism that can be realized when the distribution is effectively aligned to the target model.
Appendix C Implementation Details
For all evaluation we utilize the following hyperparameter setups. If not explicitly noted here, parameters are consistent with those specified by the authors of MDLM [18]. For reported results on CNN/DM and OWT, 200 iterations are conducted on samples randomly selected from the datasets. For evaluation on MT bench, 480 iterations are conducted.

### v4

@@ 6.1 Experimental Setup
To empirically evaluate the improvements provided by using SpecDiff, the paper provides an empirical analysis on three standard natural language processing tasks: (1) text summarization using the CNN/DM dataset Nallapati et al. (2016), (2) text generation on the OpenWebText (OWT) dataset Gokaslan et al. (2019), and (3) text generation using MT Bench Zheng et al. (2023). Additionally, we assess the performance of our method against the current state-of-the-art using SpecBench, a unified evaluation platform for speculative decoding techniques Xia et al. (2024).
In each setting, the model is queried for 1024 tokens using a greedy decoding scheme (temperature = 0). For the experiments, we evaluate the pretrained target models $M_{p}$ GPT-2 XL (1.5B), GPT-NEO (2.7B), and Vicuna (33B). We use Masked Diffusion Language Model (110M) as our drafter model $M_{q}$, which is a comparable size to the baseline drafter GPT-2 (86M) employed for our standard speculative decoding baseline Sahoo et al. (2024).
All evaluation is conducted on two NVIDIA A100 series GPUs (80GB) using CUDA 12.2. Additionally, FlashAttention Dao et al. (2022) is used to optimize the performance in all experiments.
@@
Our method is assessed empirically by walltime speedup, acceptance rate $\alpha$, and total floating point operations (FLOPs). The reported results are compared to recognized baselines of vanilla autoregressive decoding, standard speculative decoding (SpS) implementations as proposed by Leviathan et al.; Chen et al., which our method most closely resembles, and Eagle-2 Li et al. (2024a), the state-of-the-art for speculative decoding methods. Additionally, we provide comprehensive comparison to these method, as well as the current fastest-to-date speculative decoding approaches, detailing the improvements SpecDiff provides in reduction of FLOPs and memory footprint in Table 1.
@@
Table 3: Evaluation on MT Bench using SpecDiff (Ours), EAGLE, EAGLE-2, and SpS with Vicuna target models. FLOPs/step computes that floating point operations for the drafter to generate each sequence. ‘Overall Speedups’ for the baselines are measured on all Spec-Bench Xia et al. (2024) tasks. denotes results collected using Spec-Bench.
6.2 Results and Discussion
@@
[…]
@@
While previous implementations of speculative decoding rely on a common architecture between the drafter and target models Leviathan et al. (2023); Chen et al. (2023), using smaller versions of the same architecture to generate draft sequences, these experiments demonstrate a robustness to using a completely different architecture for sequence drafting. This is particularly significant given the absence of finetuning in the reported results (Tables 2 and 4). Pretrained diffusion models can be directly purposed as draft models requiring no additional training.
@@ 8 Limitations
We note that available discrete diffusion models are based on the GPT-2 tokenizer, which would restrict one immediate compatibility with target models using different tokenization schemes. For our experiments, we indeed trained a new discrete diffusion model from scratch with a different tokenizer, a process that demanded substantial computational resources and time. All our models will be released on HugginFace and thus be directly used by the community.
@@
9 Acknowledgements
This work is partially supported by NSF awards 2143706, 2232054, and 2242931. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 and was supported by the LLNL-LDRD Program under Project No. 24-ERD-010 (LLNL-CONF-2000386). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ASCR-ERCAP0027427. The views and conclusions provided in this paper reflect those of the authors only.
@@ Appendix A Drafter Convergence Analysis
In prior studies, theoretical results have focused on determining the optimal $\gamma$ to maximize the throughput of speculative decoding methods (Leviathan et al., 2023). For SpecDiff, extending $\gamma$ introduces minimal overhead and becomes less important to consider. The number of sequential diffusion operations, $T$, instead impacts $\alpha$ as increasing this number improves the convergence of $M_{q}(x)$ to the learned distribution, aims to closely approximate the distribution of $M_{p}(x)$. Hence, an implicit dependency arises between $T$ and $\alpha$, which is reflected in Figure 2 (right).
@@
We are now ready to connect this to Equation 3:
Appendix B Accepted Draft Lengths
Figure 3: Accepted draft lengths for OpenWebText evaluation.
Figure 4: Accepted draft lengths for MT Bench evaluation.
We notice a starch constrast in the lengths of accepted drafts between these two settings. While the distribution of accepted draft lengths in Figure 4 is what would be anticipated given the lower values of $\alpha$, the longer generations on OpenWebText (Figure 3) speak to the parallelism that can be realized when the distribution is effectively aligned to the target model.
Appendix C Implementation Details
For all evaluation we utilize the following hyperparameter setups. If not explicitly noted here, parameters are consistent with those specified by the authors of MDLM Sahoo et al. (2024). For reported results on CNN/DM and OWT, 200 iterations are conducted on samples randomly selected from the datasets. For evaluation on MT bench, 160 iterations are conducted.